OpenStudy (anonymous):

Jamie is 5 years older than Ella. Jamie's age is 11 years less than three times Ella's age. The system below models the relationship between Jamie's age (j) and Ella's age (e): j = e + 5 j = 3e − 11 Which of the following methods is correct to find Jamie's and Ella's age? A) solve j + 5= 3j - 11 to find the value of j B) write the points where the graph of the equation intersect the x-axis C) write the points where the graph of the equation intersect the y-axis D)

OpenStudy (anonymous):

@Mehek14

Mehek (mehek14):

what's D?

OpenStudy (anonymous):

D) solve e+5 = 3e - 11 to find the value of e.

OpenStudy (anonymous):

@Mehek14

Mehek (mehek14):

I would say D e + 5 = 3e - 11 3e - 11 = e + 5 3e - e = 5 + 11 2e = 16 e = 16/2 e = 8 j = 8 + 5 j = 13 because you find the value of e, substitute it into the equation and find j

OpenStudy (anonymous):

omgg thankssss ssooo muchh @Mehek14

Mehek (mehek14):

yourrrrrrrrrrrr welcomeeeeeeeee :))))))))